Supplemental Operation – February 25th

Supplemental Operations Report 25Feb2022

Name of person filing report: Shannon Rupert

Reason for Report: Changing of the guard

Non-nominal systems: GreenHab

Action taken for non-nominal systems: All crops were removed from the GreenHab and it will be shut down quite possibly for the rest of the field season

Generator: Still being manually run at night. 967.5 hours

ScienceDome Dual Split: Off

Solar—Charge controllers off at night. We turn the generator on before the SOC reaches 70%.

Solar— VDC Last daylight hours: (information obtained from magnum data)

Average 52.31

Minimum 48.40

Maximum 59.20

Notes on solar: Nothing to report

Propane Reading, station tank – 70 %

Propane Reading, director tank— 65 %

Propane Reading, intern tank— 55 %

Propane Reading, generator— 55 %

Ethanol Free Gasoline – 2 gallons

Water (Outpost tank) – 200 gallons

Sojourner rover used: Yes

Hours: 161.9

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): No

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, local

CrewCar used and why, where? Yes. Atila took it to drive to SLC to pick me up but lost control and went off the road close to I-70 on Highway 24. He was shaken and sore afterwards but the car was checked and appears to have no damage.

Luna used and why, where? No.

Campus wide inspection, if action taken, what and why? I have not completed my inspection fully but noted no damage/changes to the Robotic Observatory exterior, instrumentation nearby or the solar panels. Signs on the road and at the entrance have no damage/changes.

General notes and comments: I arrived back at MDRS on Wednesday night and Atila will be leaving MDRS on Saturday morning.

Summary of general operations:

Summary of internet: The original Hab internet service will be deactivated on March 1, 2022. Today the new system was to be installed, but the installer cancelled. I will be installing a Starlink system that hopefully can be used while we are waiting for HughesNet to get its act together.

Summary of suits and radios: Nothing to report

Summary of Hab operations: The toilet holding tank was cleaned for the built up waste that was making the tank fill twice as fast as when nominal. We were unable to see how successful we were, but will track the crew’s use to see if we cleared it completely. We will repeat this process in 4 weeks.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: I inspected the GreenHab the morning after I arrived. From what few details I was getting from reports and photos, it was clear something was wrong. Nothing was thriving and plants were dying and underproducting. A more detailed inspection showed an unchecked infestation of white fly and contamination of soil by eggs and larva. I determined a collapse was near, salvage wasn’t possible, and we shut the building down. The crew removed all crops and consolidated the soil, and I will be finishing the removal of plant debris over the next few days. Heat is off and supplemental lighting turned off. This failure was the result of not one, but many small things. I was offsite and hadn’t been paying enough attention to the reports to realize that the crew was not reporting vital information about the condition of the plants. Our supplemental operations report had become truncated in a way that didn’t allow clear reporting on GreenHab issues by Atila. The GreenHab Officers and Atila didn’t have the knowledge and experience to recognize what was happening and I was offsite. Although we are requiring a general skill set for GreenHab Officer’s beginning next season, we do not have any requirements at the moment and it is one of two positions that are generally not filled by qualified people. Finally, but maybe not of much importance, I requested that the crew install a shade cloth weeks ago, and that was never gone. Dying plants were reported as drying plants. This is, however, why we do what we do—to learn from what mistakes we make. We probably won’t be able to produce anymore crops this season, because we need to sterilize the soil before we can replant and that requires heat, sunlight and time.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Robotic observatory is down. A part to repair it is due here Saturday.

Summary of health and safety issues: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site